# Fast quantum integer multiplication with zero ancillas

### v3

@@ IV Space-time trade-offs
Circuit depth—The recursive tree structure of our algorithms lends itself well to parallelization, as multiple branches of the tree can be performed simultaneously (Fig. 1). Since each register is divided into $k$ parts, ostensibly it is possible to perform $k$ recursive calls in parallel without the need for ancillas. The challenge is that the $k$ linear combinations $(A\mathbf{x})_{\ell}$ must be simultaneously computed in-place. We do not know of any generic way to compute many specific linear combinations reversibly and in-place like this; instead we have designed sequences of operations by hand that are optimal or nearly optimal, which we record in Appendix F. By parallelizing in this way (and using sub-linear-depth circuits for the additions [21]), PhaseProduct can be implemented in depth $\mathcal{O}(n^{\log_{k}2})$ and PhaseTripleProduct in $\mathcal{O}(n^{\log_{k}3})$. These scalings are sublinear for $k>2$ and $k>3$ respectively. This sublinear-depth construction uses $\mathcal{O}(n)$ ancilla qubits because the sublinear depth additions require $\mathcal{O}(n)$ ancillas. We also note that although the PhaseProduct and PhaseTripleProduct can be implemented in sublinear depth, surprisingly, for the full multiplication circuit we seem to be limited by the depth of the quantum Fourier transforms; indeed, as far as we are aware, achieving a sub-linear depth QFT requires a super-linear number of ancilla qubits.
@@ V.1 Classical-quantum multiplication: Shor’s algorithm
Perhaps the most obvious application of our result is to Shor’s algorithm for integer factorization, which can be implemented via $\mathcal{O}(n)$ controlled, in-place classical-quantum modular multiplications. In particular, it is possible to implement these multiplications to precision $\eta$ using a total of $2n+\mathcal{O}(\log(1/\eta))$ qubits (see Appendix D for details). Thus, asymptotically, our technique yields a circuit for factoring with $\mathcal{O}(n^{2+\epsilon})$ gates (for arbitrarily small $\epsilon$) and only $2n+\mathcal{O}(\log n)$ qubits. To our knowledge, this is by far the best qubit count of any factoring circuit with a sub-cubic number of gates.
@@
In order to get a sense of circuit sizes in practice, we perform a detailed analysis of the gate and qubit counts for the out-of-place multiplication of 2048-bit inputs (which is used to construct the in-place multiplications described above). Specifically, we build the recursive tree (Fig. 1) and carefully tabulate the number and type of gates in each layer. Via a careful choice of the points $w_{\ell}$, we are able to devise efficient addition sequences that minimize the overhead of forming linear combinations such as $(A\mathbf{x})_{\ell}$. Moreover, at each layer, we explicitly optimize over $k$, the number of pieces into which we divide the inputs (see e.g. Sec. II.1 and Table 1). Interestingly, we find that it is sometimes optimal to use different values of $k$ even at the same level of recursion (see Appendix E).
@@ C.2 Full case
Appendix D In-place modular multiplication
@@ 4 Uncompute the ancilla qubit by computing whether $cx\bmod N<Nw$ via a comparison operator
Appendix E Estimation of gate and qubit counts
@@
E.1 Recursive structure and choice of $k$
@@ E.2 Evaluation points $w_{\ell}$
E.3 2048-bit classical-quantum multiplication
@@
E.4 $x^{2}\bmod N$ proof of quantumness
@@
Appendix F Parallel sequences
@@
Table 5: $\mathbf{k=3}$ quantum-quantum multiplication sequence. In this table only the state of the $x$ sub-registers are shown; the same operations are applied to the $z$ register. Registers containing values upon which the algorithm is applied again recursively are highlighted in bold. The linear combinations used here for the products correspond to the evaluation points $w_{\ell}\in\{0,\infty,\pm 1,\pm 2,-1/2\}$.
@@
Table 6: $\mathbf{k=4}$ quantum-quantum multiplication sequence. In this table only the state of the $x$ sub-registers are shown; the same operations are applied to the $y$ and $z$ registers. Registers containing values upon which the algorithm is applied again recursively are highlighted in bold. The linear combinations used here for the products correspond to the evaluation points $w_{\ell}\in\{0,\infty,\pm 1,\pm 1/2,\pm 2,\pm 4\}$.

### v4

@@ IV Space-time trade-offs
Circuit depth—The recursive tree structure of our algorithms lends itself well to parallelization, as multiple branches of the tree can be performed simultaneously (Fig. 1). Since each input register is divided into $k$ parts, it is possible to perform $k$ recursive calls in parallel by storing a different linear combination in each part. By parallelizing in this way, PhaseProduct can be implemented in depth $\mathcal{O}(n^{\log_{k}2})$ and PhaseTripleProduct in $\mathcal{O}(n^{\log_{k}3})$. These scalings are sublinear for $k>2$ and $k>3$ respectively, and asymptotically the exponents can be made arbitrarily close to zero by increasing $k$.
Achieving sublinear depth for PhaseProduct and PhaseTripleProduct does seem to require a small (sublinear) number of ancilla qubits, as sublinear depth addition (which is needed to compute the linear combinations $(A\bm{x})_{\ell}$) seems to require ancillas [34]. [^7] It is also not immediately clear how to parallelize the dirty qubit techniques of Appendix C into sublinear depth. However, surprisingly, the space bottleneck for sublinear-depth multiplication seems not to to be the phase rotation algorithms, but rather the quantum Fourier transforms on the output register. Indeed, to our knowledge the most space-efficient sublinear-depth (approximate) quantum Fourier transform requires $\mathcal{O}(n\log n)$ ancillas [26]. With this in mind, we formalize our result regarding the depth of quantum multiplication in the following claim, leaving the cost of the quantum Fourier transform abstract to allow for a choice of implementation:
Claim 1.
Suppose there exists a circuit for the quantum Fourier transform on $n$ qubits having depth $D_{\mathsf{QFT}}(n)$, ancilla count $A_{\mathsf{QFT}}(n)$ and gate count $G_{\mathsf{QFT}}(n)$. Then, there exist circuits for quantum integer multiplication of $n$-qubit values (Eqs. 1 and 2) with depth $\mathcal{O}(n^{\epsilon})+2D_{\mathsf{QFT}}(n)$ and ancilla count $\max(\mathcal{O}(n/\log n),A_{\mathsf{QFT}}(n))$ using $\mathcal{O}(n^{1+\epsilon})+2G_{\mathsf{QFT}}(n)$ gates, for any $\epsilon>0$.
Proof.
See Appendix D. ∎
A challenge in achieving the low depth and low qubit count of Claim 1 is that $k$ linear combinations $(A\mathbf{x})_{\ell}$ must be computed in-place, each overwriting part of the |$\ket{x}$ register—but in general each $(A\mathbf{x})_{\ell}$ depends on all the bits of |$\ket{x}$, including those being overwritten by other values! While we show in the proof of Claim 1 that it is indeed possible to perform the desired computation in-place in polylogarithmic depth, in practice this step could lead to large constant factors if done without care. Thus, in Appendix G, we propose explicit sequences of parallel operations designed by hand for a few practically-relevant values of $k$, with the $w_{\ell}$ chosen carefully so that the linear combinations can be computed in-place via a small number of additions. Finding such efficient sequences for more values of $k$ may be important if the low-depth constructions are to be used in practice.
Finally, we note that with the use of a superlinear number of qubits, we may parallelize all $q$ recursive calls into a single layer. Doing so yields an asymptotic depth of $\mathcal{O}(\log^{2}n)$, and a total qubit count of $\mathcal{O}(n^{1+\epsilon})$ for any $\epsilon>0$. It is already known that multiplication can be performed in depth $\mathcal{O}(\log^{2}n)$ using $\mathcal{O}(n\log n\log\log n)$ qubits via a parallel quantum version of the Schönhage-Strassen algorithm [35]; however, that algorithm has large constant factors, and it may be the case that for certain values of $n$ and $\epsilon$ our circuit uses fewer qubits and/or less depth in practice. The exploration of the practical costs of this extremely low-depth construction may be an interesting direction for future work.
@@ V.1 Classical-quantum multiplication: Shor’s algorithm
Perhaps the most obvious application of our result is to Shor’s algorithm for integer factorization, which can be implemented via $\mathcal{O}(n)$ controlled, in-place classical-quantum modular multiplications. In particular, it is possible to implement these multiplications to precision $\eta$ using a total of $2n+\mathcal{O}(\log(1/\eta))$ qubits (see Appendix E for details). Thus, asymptotically, our technique yields a circuit for factoring with $\mathcal{O}(n^{2+\epsilon})$ gates (for arbitrarily small $\epsilon$) and only $2n+\mathcal{O}(\log n)$ qubits. To our knowledge, this is by far the best qubit count of any factoring circuit with a sub-cubic number of gates.
@@
In order to get a sense of circuit sizes in practice, we perform a detailed analysis of the gate and qubit counts for the out-of-place multiplication of 2048-bit inputs (which is used to construct the in-place multiplications described above). Specifically, we build the recursive tree (Fig. 1) and carefully tabulate the number and type of gates in each layer. Via a careful choice of the points $w_{\ell}$, we are able to devise efficient addition sequences that minimize the overhead of forming linear combinations such as $(A\mathbf{x})_{\ell}$. Moreover, at each layer, we explicitly optimize over $k$, the number of pieces into which we divide the inputs (see e.g. Sec. II.1 and Table 1). Interestingly, we find that it is sometimes optimal to use different values of $k$ even at the same level of recursion (see Appendix F).
@@ C.2 Full case
Appendix D Proof of circuit depth
In this section of the Appendix, we prove Claim 1, originally stated in Section IV and reproduced here:
See 1
Proof.
The claim is equivalent to the statement that PhaseProduct and PhaseTripleProduct can be implemented in depth $\mathcal{O}(n^{\epsilon})$ using $\mathcal{O}(n/\log n)$ ancilla qubits and $\mathcal{O}(n^{1+\epsilon})$ gates, since these operations conjugated by the QFT implement multiplication (and the ancilla qubits can be shared between the QFTs and the phase rotation circuits). Thus we proceed by proving that statement. For simplicity, we first give the proof for PhaseProduct; at the end we show that extending the proof to PhaseTripleProduct only requires trivial modifications.
Let $\{w_{\ell}\}$ be the set of $q=2k-1$ points at which the polynomials will be evaluated. Let $A_{1}$ be the $k\times k$ Vandermonde matrix formed from the first $k$ points, $w_{0}$ through $w_{k-1}$, and let $A_{2}$ be the $(k-1)\times k$ Vandermonde matrix for the $k-1$ points $w_{k}$ through $w_{q-1}$. The broad structure of the algorithm is as follows: first, the vectors $\bm{x}$ and $\bm{z}$ are each multiplied in-place by the matrix $A_{1}$, and then PhaseProduct is called in parallel on each of the $k$ elements of the result. Once those recursive calls are complete, the matrix-vector product with $A_{1}$ is inverted, and then the inputs are multiplied by $A_{2}$, such that PhaseProduct can again be applied in parallel, for the remaining $k-1$ $w_{\ell}$. Finally the matrix-vector product with $A_{2}$ is inverted, returning the vectors to their original values.
[…]
  • Depth: $D(t)=2D(t/k)+2D_{\mathcal{C}_{A_{1}}}(t)+2D_{\mathcal{C}_{A_{2}}}(t)$
  • Ancillas: $A(t)=kA(t/k)+2\cdot\max(S_{\mathcal{C}_{A_{1}}}(t),S_{\mathcal{C}_{A_{2}}}(t))$
  • Gates: $G(t)=(2k-1)G(t/k)+4G_{\mathcal{C}_{A_{1}}}(t)+4G_{\mathcal{C}_{A_{2}}}(t)$
In the next paragraphs we show how to construct $\mathcal{C}_{A_{1}}$ and $\mathcal{C}_{A_{2}}$ efficiently, in order to achieve the complexities stated in the Claim.
The in-place matrix-vector product is straightforward for triangular matrices, by iterating through the matrix one row at a time (starting with the top row for an upper triangular matrix, and the bottom row for a lower triangular one). The matrices $A_{1}$ and $A_{2}$ are full rank, and thus can each be decomposed into the product of two triangular matrices via the LU decomposition (up to reordering of the rows), which should be precomputed classically. As shown in Appendix B, one may arbitrarily rescale rows of $A_{1}$ and $A_{2}$, which can be used to ensure that the LU decomposition consists entirely of integer values. Thus the quantum in-place matrix-vector product can thus be computed simply via a series of integer products, some in-place (corresponding to diagonal elements of the LU matrices) and some out-of-place (corresponding to off-diagonal elements). The total number of integer products is proportional to the size of the matrices $A_{1}$ and $A_{2}$, which is independent of $n$. For all of the integer products, one input is quantum and the other is an $O(1)$-bit classical value.
[…]
Thus, we find that overall, the matrix-vector multiplications can be performed in $D_{\mathcal{C}_{A}}(t)=\mathcal{O}(\operatorname{polylog}(t))$, $S_{\mathcal{C}_{A}}(t)=\mathcal{O}(t/\log^{2}t)$, and $G_{\mathcal{C}_{A}}(t)=\mathcal{O}(t)$. Using these values in the recursion relations for PhaseProduct yields the complexities:
  • Depth: $D(n)=\mathcal{O}(n^{\log_{k}2})$
  • Ancillas: $A(n)=\mathcal{O}(n/\log n)$
  • Gates: $G(n)=\mathcal{O}(n^{\log_{k}(2k-1)})$
The choice of when to stop the recursion is important to achieve this ancilla qubit count. If we were to recurse to a base case of size $O(1)$, each of the $O(n)$ leaves of the recursive tree would require $O(1)$ ancillas. Instead we observe that a base case of size $t_{b}$ can be implemented via the schoolbook algorithm in depth $O(t_{b})$ using zero ancillas; thus, we can switch to the base case at a size of $t_{b}=\mathcal{O}(n^{\log_{k}2})$ without adversely affecting the depth, which maintains the sublinear qubit count.
The same proof applies to PhaseTripleProduct, except that $q=3k-2$, and we thus must apply the recursive calls in three layers, instead of two. This yields the following complexities for PhaseTripleProduct:
  • Depth: $D(n)=\mathcal{O}(n^{\log_{k}3})$
  • Ancillas: $A(n)=\mathcal{O}(n/\log n)$
  • Gates: $G(n)=\mathcal{O}(n^{\log_{k}(3k-2)})$
For both PhaseProduct and PhaseTripleProduct, for any $\epsilon>0$ there exists a value of the constant $k$ which yields depth smaller than $O(n^{\epsilon})$ and gate count smaller than $O(n^{1+\epsilon})$, completing the proof. ∎
Lemma 1.
Given an $n$-qubit register |$\ket{x}$ and $O(1)$-bit classical constant $c$, the isometry |$\ket{x}\to\ket{cx}\ket{g(x)}$, where $g$ is unspecified “garbage” data, can be performed in depth $\mathcal{O}(\operatorname{polylog}(n))$ and gate count $O(n)$ using a total of $n+O(n/\log^{2}n)$ qubits including the input register.
Proof.
The folowing algorithm achieves the desired operation:
  1. 1.
    Let $d$ be the integer such that $c^{\prime}=c/2^{d}$ is an odd integer.
  2. 2.
    Divide |$\ket{x}$ into blocks of $m=O(\log^{2}n)$ qubits each. Label the value of block $i$ |$\ket{X_{i}}$.
  3. 3.
[…]
  4. 4.
    Increment the entire input register by $b=\sum_{i}2^{mi}B_{i}$.
  5. 5.
    Multiply the input register by $2^{d}$ (via a bit shift).
We now briefly explain each step, and how it can be performed efficiently. Steps 1 and 2 are entirely classical. In step 3, the mapping of |$\ket{X_{i}}\to\ket{c^{\prime}X_{i}}$ on each block (here not distinguishing between overflow and input qubits) can be implemented in depth $O(\operatorname{polylog}n)$ using $O(1)$ ancillas per block, by performing an in-place addition of the value $2^{j}(c^{\prime}-1)$ controlled off of each bit $j$ of $X_{i}$. By iterating from the most significant bit of $X_{i}$ to the least significant, it is ensured that later control bits are not affected by previous additions. Here we may use standard linear-depth adders, because the size of the additions is the block size $m=\mathcal{O}(\log^{2}n)$. At the end of step 3, the full state can be written |$\ket{c^{\prime}x-b}\ket{b}$ for $b=\sum_{i}2^{mi}B_{i}$ as above, where here the first register is the input register, and the second is a register of $O(n/\log^{2}n)$ ancilla qubits representing $b$ in a sparse format of $O(1)$ qubits per $m$-qubit block (the rest of the bits of $b$ are zero). Step 4 then maps the state to |$\ket{c^{\prime}x}\ket{b}$; this mapping can be performed in depth $\mathcal{O}(\log n)$ using $O(1)$ ancillas per block (that are returned to |$\ket{0}$ immediately) via a low-space quantum carry-lookahead adder construction [34], modified slightly to account for the fact that $b$ is sparse. Finally, the bitshift of step 5 by a classical constant $d$ yields a final state |$\ket{cx}\ket{b}$, which is the desired final state (with $g(x)=b$). ∎
Appendix E In-place modular multiplication
@@ 4 Uncompute the ancilla qubit by computing whether $cx\bmod N<Nw$ via a comparison operator
Appendix F Estimation of gate and qubit counts
@@
F.1 Recursive structure and choice of $k$
@@ F.2 Evaluation points $w_{\ell}$
F.3 2048-bit classical-quantum multiplication
@@
F.4 $x^{2}\bmod N$ proof of quantumness
@@
Appendix G Parallel sequences
@@
Table 5: $\mathbf{k=3}$ parallel sequence for PhaseTripleProduct. In this table only the state of the $x$ sub-registers are shown; the same operations are applied to the $y$ and $z$ registers. Registers containing values upon which the algorithm is applied again recursively are highlighted in bold. The linear combinations used here for the products correspond to the evaluation points $w_{\ell}\in\{0,\infty,\pm 1,\pm 2,-1/2\}$.
@@
Table 6: $\mathbf{k=4}$ parallel sequence for PhaseTripleProduct. In this table only the state of the $x$ sub-registers are shown; the same operations are applied to the $y$ and $z$ registers. Registers containing values upon which the algorithm is applied again recursively are highlighted in bold. The linear combinations used here for the products correspond to the evaluation points $w_{\ell}\in\{0,\infty,\pm 1,\pm 1/2,\pm 2,\pm 4\}$. We note that it may be possible to parallelize the products further into three layers, rather than four, but the number of required additions would likely increase.